# IV. Characterizing atmosphere gray extinction with thermal infrared observations

### v1

@@ C.4 Mie scattering of aerosols
Figure 13: Synthetic transmission curve computed with libRadtran at airmass $X$ = 1 for the OHP site location (solid black line) with the same atmosphere as Fig. 1. The Earth’s atmosphere attenuates the flux from stellar sources through Rayleigh scattering (purple dotted curve), Mie scattering by aerosols with $\mathrm{\tau_{0}^{aer}}$ = 0.05 and $\mathrm{\beta^{aer}}$ = 1.4 (green dashed curve), molecular absorption by oxygen $\mathrm{O_{2}}$ (red dashed-dotted curve), ozone $\mathrm{O_{3}}$ = 350 DU (blue wide-dashed curve), and water vapour PWV = 12 mm (light blue dotted curve). The instrumental response of the StarDICE telescope ugrizy photometric system is also shown for reference (Souverin et al. 2025).
@@ Appendix D Impact of stellar parameters and instrumental throughput errors on chromatic extinction estimation
Both the sources stellar parameters from the Gaia DR3 catalog and the instrumental transmission come with uncertainties which propagate onto the synthetic spectra and the total throughput respectively. We do have an estimation on stellar parameters being $\sim$180 K on $\mathrm{T_{eff}}$ and 0.2 dex on $\log\mathrm{g}$ and [Fe/H] (Fouesneau et al. 2023). The propagated error on $\mathcal{C}_{s,i}$ for the stellar parameters uncertainty is less than 0.2 mmag for a star with $\mathrm{T_{eff}}$ = 5000 K and $\log\mathrm{g}=0.5$ and $\mathrm{[Fe/H]}=0$. It is difficult to estimate the uncertainty on instrumental transmissions for which we use typical transmissions reported by manufacturer’s datasheets. The shift on $\mathcal{C}_{s,i}$ is approximately 1 mmag when adopting a conservative slope error of 30% on the shape of the total instrumental throughput curve. Overall, these errors are negligible compared to the photometric measurements uncertainties which are at least one order of magnitude higher.

### v2

@@ C.4 Mie scattering of aerosols
Figure 13: Synthetic transmission curve computed with libRadTran at airmass $X$ = 1 for the OHP site location (solid black line) with the same atmosphere as Fig. 1. The Earth’s atmosphere attenuates the flux from stellar sources through Rayleigh scattering (purple dotted curve), Mie scattering by aerosols with $\mathrm{\tau_{0}^{aer}}$ = 0.05 and $\mathrm{\beta^{aer}}$ = 1.4 (green dashed curve), molecular absorption by oxygen $\mathrm{O_{2}}$ (red dashed-dotted curve), ozone $\mathrm{O_{3}}$ = 350 DU (blue wide-dashed curve), and water vapour PWV = 12 mm (light blue dotted curve). The instrumental response of the StarDICE telescope ugrizy photometric system is also shown for reference (Souverin et al., 2025).
@@ Appendix D Impact of stellar parameters and instrumental throughput errors on chromatic extinction estimation
Both the sources stellar parameters from the Gaia DR3 catalog and the instrumental transmission come with uncertainties which propagate onto the synthetic spectra and the total throughput respectively. We do have an estimation on stellar parameters being $\sim$180 K on $\mathrm{T_{eff}}$ and 0.2 dex on $\log\mathrm{g}$ and [Fe/H] (Fouesneau et al., 2023). The propagated error on $\mathcal{C}_{s,i}$ for the stellar parameters uncertainty is less than 0.2 mmag for a star with $\mathrm{T_{eff}}$ = 5000 K and $\log\mathrm{g}=0.5$ and $\mathrm{[Fe/H]}=0$. It is difficult to estimate the uncertainty on instrumental transmissions for which we use typical transmissions reported by manufacturer’s datasheets. The shift on $\mathcal{C}_{s,i}$ is approximately 1 mmag when adopting a conservative slope error of 30 per cent on the shape of the total instrumental throughput curve. Overall, these errors are negligible compared to the photometric measurements uncertainties which are at least one order of magnitude higher.
[…]